OpenStudy (anonymous):

An oxygen tank contains oxygen at a pressure of 70 kPa. When the pressure drops to 36 kPa, the temperature increases from 300 K to 350 K. If the initial volume is 9 L, what is the final volume? --------------- 20.4 L 84.0 L 233 L 336 L

OpenStudy (anonymous):

again you begin with: pv=nrt n=n n=pV/RT p1V1/RT1=p2V2/RT2 p1V1/T1=p2V2/T2 p1V1T2=p2V2T1 V2 = p1V1T2/p2T1

OpenStudy (anonymous):

@Austin_Cochran HELP!

OpenStudy (anonymous):

you just need to put the numbers in equation i gave you and you get that volume is: 20.4 L

OpenStudy (anonymous):

oh! thank you , you're just too smart for me